OpenStudy (karlaltr):

Activity of technetium-99 in 24 hours Time (h) 0 3 6 9 12 15 18 21 24 Acivity (kBq) 17.0 12.2 8.9 6.5 4.5 3.2 2.3 1.5 1.1 determine the half life of technetium-99

OpenStudy (karlaltr):

HEeres a better picture of the graph

OpenStudy (karlaltr):

i will fan and medal PLEASE HELP

OpenStudy (anonymous):

what music u listen to

OpenStudy (karlaltr):

can you help me with this question? :p but i listen to hiphop and rap....anytime of music is what i listen to tbh

OpenStudy (dariusx):

T=ln2/λ=0.693/λ when t = 0, N = No n = m/M M = 99g/mol N = n * Na Na = 6,02 * 10^23, A=λ∗N idk, how to solve it fully, but imma throw some formulas here. maybe someone will find them useful and will help you. if i knew 'm' i could solve it

OpenStudy (karlaltr):

its really confusing since we dont have m, theres now way of finding it!

OpenStudy (karlaltr):

but thank you so much!

OpenStudy (karlaltr):

Idk if this wil help but the formula My teacher provided was A=A0(1/2) t/h